9.The perimeter of a rectangular farm is 1800 m and its length is 140 m longer than its width. Find the area of the farm. (in simultaneous equations)
I've already established that L = W + 140 and that LxW= 1800.
thanks :)

Answer by mananth(16946)

perimeter = 2(L+W)=1800

L+W=900
L-w=140
Find L & W and multiply to get area